Find the perimeter of a rectangle whose length is 5 cm less than twice its width. Its width is 15cm
Answer by checkley77(12569)

L=2W-5
W=15
L=2*15-5
L=30-5
L=25 ANS.
P=2L=2W
P=2*25+2*15
P=50+30
P=80 ANS. FOR THE PERIMETER.